# Long-horizon Locomotion and Manipulation on a Quadrupedal Robot with Large Language Models

### v2

@@ Abstract
We present a large language model (LLM) based system to empower quadrupedal robots with problem-solving abilities for long-horizon tasks beyond short-term motions. Long-horizon tasks for quadrupeds are challenging since they require both a high-level understanding of the semantics of the problem for task planning and a broad range of locomotion and manipulation skills to interact with the environment. Our system builds a high-level reasoning layer with large language models, which generates hybrid discrete-continuous plans as robot code from task descriptions. It comprises multiple LLM agents: a semantic planner for sketching a plan, a parameter calculator for predicting arguments in the plan, and a code generator to convert the plan into executable robot code. At the low level, we adopt reinforcement learning to train a set of motion planning and control skills to unleash the flexibility of quadrupeds for rich environment interactions. Our system is tested on long-horizon tasks that are infeasible to complete with one single skill. Simulation and real-world experiments show that it successfully figures out multi-step strategies and demonstrates non-trivial behaviors, including building tools or notifying a human for help. Demos are available on our project page: https://sites.google.com/view/long-horizon-robot.
@@ I Introduction
Figure 1: Overview of the hierarchical system for long-horizon loco-manipulation task. The system is built up from a reasoning layer for task decomposition and a controlling layer for skill execution. Given the language description of the long-horizon task, a cascade of LLM agents perform high-level task planning and generate function calls of parameterized robot skills. The controlling layer instantiates the mid-level motion planning and low-level controlling skills with RL.
We propose a hierarchical system to address the challenges in long-horizon loco-manipulation problems for quadrupeds as illustrated in Fig. 1. At the high level, we harness the planning ability of pretrained large language models (LLMs) for abstract reasoning over the long horizon. Leveraging the wealth of encoded common-sense knowledge, the LLM-based reasoning layer can directly predict high-level strategies without demonstrations or searching heuristics. To better ground to robotic problems, we design the reasoning module as a cascade of multiple LLM agents. A semantic planner sketches a discrete plan, while a parameter calculator predicts continuous parameters. Subsequently, a code generator encapsulates the plan into function calls for robot skills. At the low level, we train a spectrum of motion planning and control policies with model-free reinforcement learning, spanning from quadrupedal locomotion to bipedal manipulation.
Our contributions are as follows: We introduce a new reasoning architecture with multiple LLM agents which focuses on long-sequence robotic problems. Our system chains together both locomotion and manipulation skills, and enables more complex tasks that require nuanced physical interaction. Combined with the high-level reasoning capabilities of LLM, it can address complex long-horizon tasks that could span over 10 minutes of execution. We extensively validate our design components in four novel and challenging long-horizon tasks for quadrupedal robots including manipulating a light switch at a height, package delivery, bridge building, and taking an elevator. Our method outperforms all the baselines, demonstrating a success rate of over 70% in simulation and successful deployment in the real world.
@@ II Related Work
Quadrupedal locomotion and manipulation: Developing versatile motions over quadrupedal robots to make them work like real animals has attracted much research interest. A large body of work focuses on robust and agile locomotion skills that can walk on challenging terrains [9, 3, 10] and perform extreme parkour [11, 12]. Some recent works develop manipulation skills for quadrupeds to better interact with the environment, such as learning manipulation via locomotion to push heavy objects [13], manipulating with one or two legs while balancing on other legs [14, 7, 6], or manipulating with a gripper mounted on the robot [15, 16]. There are also works that learn mid-level motion planning policies over these low-level motion controllers for more complex skills, such as soccer manipulation [17, 18]. We focus on empowering quadrupeds with long-horizon problem-solving abilities beyond motion planning and control skills for specific tasks. The problems considered in this work require an abstract task-planning ability that reasons over the semantics. The desired strategy, such as building stairs alongside the wall so that the robot can climb up and reach the light switch, requires common-sense knowledge about the physical constraints of the environment and the robot’s capabilities and is challenging to discover from scratch.
@@
The most relevant work to us is RoboTool [33] which also reasons over long-range physical puzzles with multiple LLMs. However, RoboTool only leverages the basic quadrupedal locomotion skills, which essentially treats the quadruped robot as a mobile platform and does not fully utilize its high dimensionality. Furthermore, RoboTool is designed to generate plans specific to certain scene configurations, while our method can generate branching conditions to handle various scene configurations with the same piece of code. As we will see in Sec. IV-B, our method shows more robust planning performance.
@@ III Method
Figure 2: Illustration of the LLM-based high-level reasoning layer that generates hybrid discrete-continuous plans from the task description in language. It is composed of a semantic planner that proposes a solution consisting of branches conditioned on environment specifications and primitive actions, a parameter calculator that fills in arguments for the actions, a code generator to summarize the plan as executable robot code. The texts are abbreviated generated contents from the LLMs.
@@ III-A High-level reasoning for long-horizon tasks with LLMs
To address the challenging reasoning and grounding problem, we design a cascade of three LLM agents with different focuses as the high-level reasoning module as shown in Fig. 2. A semantic planner focuses on task decomposition with constraints from physical affordances and robot capabilities and sketches a plan of primitive skills for the long-horizon problem. A parameter calculator figures out the precise arguments for the primitives in the plan. A code generator converts the plan and the arguments into one single piece of executable code that can strategically combine the available low-level skills to handle the long-horizon problem. We use GPT-4-turbo-preview [34] for all the LLM agents.
@@ III-B2 Learning motion planning policies with hierarchical RL
III-B3 Chaining RL skills
Directly executing different skills sequentially often leads to suboptimal behavior since the terminal state distribution of one policy may not perfectly match the initial state distribution of the next policy.
For instance, the preceding box-pushing policy often terminates with the box deviating from the ideal pose right beneath the existing stairs. Executing the climbing skill from the out-of-distribution states will lead to missteps and falls, thus resulting in lower success rates as more skills are executed. To mitigate the impact of the preceding strategy’s termination state on the subsequent strategy, we collect a set of termination states of the preceding policy of each policy, then fine-tune the policy from a mixture of the original initial state distribution and the collected states. Note that we only run fine-tuning for the pair of policies that will be executed sequentially according to the plan generated by the reasoning layer to reduce the computation burden.
@@ IV-A Simulation experiment setting
Figure 4: Total number of occurrences for different error types in generated code, with ten samples per variant. Spatial error refers to errors in perceiving spatial relations leading to miscalculations, constraint violation refers to ignoring the given constraints, and logical error refers to faulty logical reasoning. The lower part shows the execution of the 4 tasks in simulation.
@@ IV-A1 Baselines
We first verify the proposed design of the high-level reasoning layer composed of a cascade of LLM agents in the simulation. We compare our method LLM (S+P+C) that leverages a semantic planner, a parameter calculator and a code generator with the following variants:
@@ IV-A3 Evaluation metrics
IV-B Ablation study analysis
As reported in Table I, removing the parameter calculator and the semantic planner leads to a significant performance drop. Note that LLM (S+C) and LLM (C) fail to achieve any success in most tasks since they mostly get stuck at the first executed skill with incorrect arguments. LLM (S+C) struggles to correctly consider the bounding box of objects and their geometric relationship. As shown in Fig. 4, the number of spatial miscalculations increases dramatically after the calculator is removed. For a more specific example, it sets the target position of climbing to the center of a stair rather than the top surface, and does not consider the robot’s height. It also decides to push the movable box exactly beneath the lower stair, which is physically infeasible. Without a semantic planner for preliminary analysis and planning, LLM (C) is more likely to make incorrect inferences and may even violate constraints, leading to ineffective or dangerous actions. In the bridge building task, it generated an incorrect execution sequence by attempting to push the box after climbing the low platform.
According to Table I, RoboTool demonstrates subpar performance across most tasks, which we attribute to three main reasons. The code generated by RoboTool lacks branching structures, limiting its adaptability to complex tasks. For instance, elevator ride requires determining whether to press the up or down button based on the current and target floors. Faced with a variety of possible combinations, RoboTool regenerates code for each new scenario, leading to inefficiency and instability. In contrast, our generated code uses an elegant branching structure that covers all possible scenarios for this task.ts. RoboTool is more likely to make mistakes in spatial calculation since it requires numerical values in the scene object description while simultaneously using variables in the prompts. As shown in Fig. 4, this mixed usage often introduced instability, leading to numerous errors in spatial relationship calculations. Furthermore, RoboTool summarizes the previous LLM agent output before passing it to the next. We found that these summaries often omit details necessary for effective task execution. Relevant prompts and code can be found from our website.
@@ IV-C Real world experiments
We test our system on two aforementioned benchmark tasks in the real world. As shown in the upper part of Fig. 5, we execute the generated code together with the RL motion skills to turn off the light. In the initial setup, the height of the light button is out of reach with any single skill, and the first level of stairs is too high for the robot to climb directly. Considering the scene configuration, the LLM planner decides to push a lower box in front of the existing stairs so that the robot can use it as an intermediate platform to climb up then to touch the button. Finally, it runs a sit-down policy to transit to the quadrupedal pose. We also show how the quadrupedal robot delivers a package into a room with the door closed initially. To get access to the room, our reasoning layer figures out a strategy to ring a bell to notify the human inside the room to open the door. To protect the package from unnecessary collisions, it leverages the obstacle-avoidance locomotion skill when moving to the bell and before pushing the package.
Failure case: The robot may enter dead ends when the execution of some low-level skills fails. Since our system gives plans in an open-loop manner, the robot cannot get an updated plan in such cases. In addition, our system cannot invent policies beyond the fixed set of low-level skills, and thus cannot recover from fatal states without a reset policy.
@@ V Conclusion
In summary, we present an LLM-based system for tackling long-horizon tasks with a quadruped robot that extends the autonomy of quadrupeds from shorter-term motions for specific skills to long-range complex behaviors that require orchestrating multiple locomotion and manipulation policies. The high-level task planning is enabled with our design of multiple LLM agents for decomposing the whole task into robot actions and then grounded to parameterized robot API calls. The low-level control that exploits the agility of quadrupeds to unlock rich environment interactions is instantiated with model-free RL. Our LLM reasoning layer plans in an open loop. Allowing it to improve from feedback such as execution traces with in-context learning is an interesting future direction. Currently, the reasoning layer plans over a fixed set of low-level robot skills crafted by experts. Based on this work, exciting future research could investigate the possibility of actively discovering new skills to learn when faced with unprecedented circumstances.

### v3

@@ Abstract
We present a large language model (LLM) based system to empower quadrupedal robots with problem-solving abilities for long-horizon tasks beyond short-term motions. Long-horizon tasks for quadrupeds are challenging since they require both a high-level understanding of the semantics of the problem for task planning and a broad range of locomotion and manipulation skills to interact with the environment. Our system builds a high-level reasoning layer with large language models, which generates hybrid discrete-continuous plans as robot code from task descriptions. It comprises multiple LLM agents: a semantic planner that sketches a plan, a parameter calculator that predicts arguments in the plan, a code generator that converts the plan into executable robot code, and a replanner that handles execution failures or human interventions. At the low level, we adopt reinforcement learning to train a set of motion planning and control skills to unleash the flexibility of quadrupeds for rich environment interactions. Our system is tested on long-horizon tasks that are infeasible to complete with one single skill. Simulation and real-world experiments show that it successfully figures out multi-step strategies and demonstrates non-trivial behaviors, including building tools or notifying a human for help. Demos are available on our project page: https://sites.google.com/view/long-horizon-robot.
@@ I Introduction
Figure 1: Overview of the hierarchical system for long-horizon loco-manipulation task. The system is built up from a reasoning layer for task planning and a controlling layer for skill execution. Given the language description of the long-horizon task, three cascaded LLM agents perform high-level task decomposition and generate a complete plan. A replanner is introduced to handle unexpected situations. The reasoning layer generates function calls for parameterized robot skills, and the controlling layer instantiates the mid-level motion planning and low-level controlling skills with RL.
We propose a hierarchical system to address the challenges in long-horizon loco-manipulation problems for quadrupeds as illustrated in Fig. 1. At the high level, we harness the planning ability of pretrained large language models (LLMs) for abstract reasoning over the long horizon. Leveraging the wealth of encoded common-sense knowledge, the LLM-based reasoning layer can directly predict high-level strategies without demonstrations or searching heuristics. To better ground to robotic problems, we design the reasoning module as a cascade of multiple LLM agents. A semantic planner sketches a discrete plan, while a parameter calculator predicts continuous parameters. Subsequently, a code generator encapsulates the plan into function calls for robot skills. To achieve high-level closed-loop planning, we introduce a replanner to handle unexpected events. At the low level, we train a spectrum of motion planning and control policies with model-free reinforcement learning, spanning from quadrupedal locomotion to bipedal manipulation.
Our contributions are as follows: We introduce a new closed-loop reasoning architecture with multiple LLM agents, which focuses on long-sequence robotic problems. It is capable of generating comprehensive plans adaptable to diverse scenes and dynamically adjusting through feedback. Our system chains both locomotion and manipulation skills, and enables more complex tasks that require nuanced physical interaction. With the high-level reasoning capabilities of LLM, it can address complex long-horizon tasks involving more than ten function calls of skills. We extensively validate our design components in four novel and challenging long-horizon tasks for quadrupedal robots including manipulating a light switch at a height, package delivery, bridge building, and taking an elevator. Our method outperforms all the baselines, demonstrating a success rate of over 70% in simulation and successful deployment in the real world.
@@ II Related Work
Quadrupedal locomotion and manipulation: Developing versatile motions over quadrupedal robots to make them work like real animals has attracted much research interest. A lot of work focuses on robust and agile locomotion skills that can walk on challenging terrains [9, 3, 10] and perform extreme parkour [11, 12]. Some recent works develop manipulation skills for quadrupeds to better interact with the environment, such as learning manipulation via locomotion to push heavy objects [13], manipulating with one or two legs while balancing on other legs [14, 7, 6], or manipulating with a gripper mounted on the robot [15, 16]. There are also works that learn mid-level motion planning policies over these low-level motion controllers for more complex skills, such as soccer manipulation [17, 18]. We focus on empowering quadrupeds with long-horizon problem-solving abilities beyond motion planning and control skills for specific tasks. The problems considered in this work require an abstract task-planning ability that reasons over the semantics. The desired strategy, such as building stairs alongside the wall so that the robot can climb up and reach the light switch, requires common-sense knowledge about the physical constraints of the environment and the robot’s capabilities and is challenging to discover from scratch.
@@
The most relevant work to us is RoboTool [33] which also reasons over long-range physical puzzles with multiple LLMs. However, RoboTool’s open-loop planning produces fixed plans for specific scene configurations, limiting its adaptability to complex tasks and unexpected events. In contrast, our approach generates branching conditions within the same code to handle diverse configurations and incorporates a replanner that dynamically adjusts the plan based on execution feedback. Additionally, RoboTool only leverages the basic quadrupedal locomotion skills, which essentially treats the quadruped robot as a mobile platform and does not fully utilize its high dimensionality. As we will see in Sec. IV-B, our method shows more robust planning performance.
@@ III Method
Figure 2: Illustration of the LLM-based high-level reasoning layer that generates primary hybrid discrete-continuous plans from the task description in language. It is composed of a semantic planner that proposes a solution consisting of branches conditioned on environment specifications and primitive actions, a parameter calculator that fills in arguments for the actions, a code generator to summarize the plan as executable robot code. The texts are abbreviated generated contents from the LLMs.
@@ III-A High-level reasoning for long-horizon tasks with LLMs
To address the challenging reasoning and grounding problem, we design four LLM agents with different focuses as the high-level reasoning module. The first three cascaded agents, semantic planner, parameter calculator, and code generator, work together to generate a complete initial plan as shown in Fig. 2. The semantic planner focuses on task decomposition with constraints from physical affordances and robot capabilities and sketches a plan of primitive skills for the long-horizon problem. The parameter calculator figures out the precise arguments for the primitives in the plan. The code generator converts the plan and the arguments into one single piece of executable code that can strategically combine the available low-level skills to handle the long-horizon problem. In addition, a replanner is introduced as the fourth agent to enable closed-loop control, addressing unexpected situations by adjusting the plan and recalculating necessary parameters to generate a new code plan dynamically during execution
We use GPT-4-turbo-preview [34] for all the LLM agents.
@@ III-A3 Code generator
III-A4 Replanner
We implement a replanner to handle unexpected events during execution, such as human interruptions that change the task goal or skill execution failures. Open-loop planning lacks the ability to adapt to execution feedback, making it vulnerable to systemic errors and unforeseen events. To mitigate these, when an unexpected event occurs, the robot first invokes a recovery strategy and gathers relevant error information, including the spatial information about the robot and the associated object of the failed skill. In the case of a human interruption, the robot receives new language prompt instructions to adapt its behavior accordingly. The replanner then leverages the original environment description and the previously generated primary code plan to produce a new plan adapted to the current situation. By extracting necessary knowledge from the primary code plan, it bypasses the first three cascaded structure, reducing the overhead of plan adjustments. By analyzing execution errors, the replanner determines whether to refine the plan or recalibrate parameters. This closed-loop approach allows the robot to iteratively correct its actions, enhancing adaptability and ensuring robust task execution.
@@ III-B2 Learning motion planning policies with hierarchical RL
Figure 4: Total number of occurrences for different error types in generated primary plan code, with ten samples per variant. Spatial error refers to errors in perceiving spatial relations leading to miscalculations, constraint violation refers to ignoring the given constraints, and logical error refers to faulty logical reasoning. The lower part shows the execution of the 4 tasks in simulation.
III-B3 Chaining RL skills and recovery
Directly executing different skills sequentially often leads to suboptimal behavior since the terminal state distribution of one policy may not perfectly match the initial state distribution of the next policy. For instance, the preceding box-pushing policy often terminates with the box deviating from the ideal pose right beneath the existing stairs. Executing the climbing skill from the out-of-distribution states will lead to missteps and falls, thus resulting in lower success rates as more skills are executed. To mitigate the impact of the preceding strategy’s termination state on the subsequent strategy, we collect a set of termination states of the preceding policy of each policy, then fine-tune the policy from a mixture of the original initial state distribution and the collected states. Note that we only run fine-tuning for the pair of policies that will be executed sequentially according to the plan generated by the reasoning layer to reduce the computation burden.
We also train an RL-based recovery policy to restore a stable quadrupedal stance. By sampling states during execution, especially in risky skills like climbing and hand touching, we define diverse initial conditions. Rewards then guide smooth transitions to target joint states, enabling rapid failure recovery and ensuring readiness for replanning.
@@ IV-A1 Baselines
We first verify the proposed design of the high-level reasoning layer composed of a cascade of LLM agents in the simulation. We present our method LLM (S+P+C+R), which incorporates a semantic planner, a parameter calculator, a code generator, and a replanner. We impose a time limit for task completion, within which the replanner can be called if a skill execution fails, and then compare it with the following variants.
  • LLM (S+P+C) does not allow the invocation of replanner but retains a full primary plan generation structure.
@@
[…]
IV-A2 Benchmark tasks
IV-A3 Evaluation metrics
IV-B Ablation study analysis
As reported in Table I, LLM (S+P+C+R) achieves the best performance across tasks. Compared to LLM (S+P+C), the closed-loop planning control provided by the replanner enables the quadruped robot to replan and retry based on error information after failing a challenging subtask (e.g., button pressing), significantly improving the overall task execution success rate. In the light task, if the robot fails to directly climb the lower stair as planned, it recognizes the system’s estimation error and replans to decisively push a box as a step, ensuring a more stable and reliable approach to success. However, for the NormDist metric, the closed-loop LLM (S+P+C+R) may retry early steps that do not lower the metric, which leads to a suboptimal score. Removing the parameter calculator and the semantic planner leads to a significant performance drop.
Note that LLM (S+C) and LLM (C) fail to achieve any success in most tasks since they mostly get stuck at the first executed skill with incorrect arguments. LLM (S+C) struggles to correctly consider the bounding box of objects and their geometric relationship. As shown in Fig. 4, the number of spatial miscalculations increases dramatically after the calculator is removed. For a more specific example, it sets the target position of climbing to the center of a stair rather than the top surface, and does not consider the robot’s height. It also decides to push the movable box exactly beneath the lower stair, which is physically infeasible. Without a semantic planner for preliminary analysis and planning, LLM (C) is more likely to make incorrect inferences and may even violate constraints, leading to ineffective or dangerous actions. In the bridge building task, it generated an incorrect execution sequence by attempting to push the box after climbing the low platform.
Figure 5: Replan deployment. a) Replan on execution failure: After a strategy fails, the robot uses a language-conditioned error detection pipeline to gather spatial information and adjust the parameters. b) Replan on interruption: When the human is unable to open the door but allows the robot to operate the door switch, the robot interrupts the original plan and presses the button to open the door, generating a new code plan while still achieving the long-term delivery goal.
@@ IV-B Ablation study analysis
Comparison with RoboTool: According to Table I, RoboTool demonstrates subpar performance across most tasks, which we attribute to three main reasons. The code generated by RoboTool lacks branching structures, limiting its adaptability to complex tasks. For instance, elevator ride requires determining whether to press the up or down button based on the current and target floors. Faced with a variety of possible combinations, RoboTool regenerates code for each new scenario, leading to inefficiency and instability. In contrast, our generated code uses an elegant branching structure that covers all possible scenarios for this task. RoboTool is more likely to make mistakes in spatial calculation since it requires numerical values in the scene object description while simultaneously using variables in the prompts. As shown in Fig. 4, this mixed usage often introduced instability, leading to numerous errors in spatial relationship calculations. Furthermore, RoboTool is an open-loop planning framework without a replanner, making it incapable of adjusting its strategy based on failure feedback. Relevant prompts and code can be found from our website.
@@ IV-C Real world experiments
Replanner Deployment: To support the real-world deployment of replan-on-failure with the necessary failure-related data, we developed a pipeline using off-the-shelf models, as shown in Fig. 5. A third-person-view camera records the robot’s policy execution. Upon failure, the robot identifies key targets for query based on the failed skill. The object detector [41] guided by the corresponding prompt predicts target bounding boxes, which are then fed into the SAM2 tracker [42] to capture the targets’ spatial trajectories during execution. Finally, we derive the geometric trajectory by comparing depth maps. For the hand touching policy, this trajectory enables the determination of the minimum distance between the robot’s toe and the button, allowing the replanner to use this offset to refine the code plan.
Long-horizon Task: We test our system on two aforementioned benchmark tasks in the real world. As shown in the upper part of Fig. 6, we execute the generated code together with the RL motion skills to turn off the light. In the initial setup, the height of the light button is out of reach with any single skill, and the first level of stairs is too high for the robot to climb directly. Considering the scene configuration, the LLM planner decides to push a lower box in front of the existing stairs so that the robot can use it as an intermediate platform to climb up then to touch the button. Finally, it runs a sit-down policy to transit to the quadrupedal pose.
We also show how the quadrupedal robot delivers a package into a room with the door closed initially. To get access to the room, our reasoning layer figures out a strategy to ring a bell to notify the human inside the room to open the door. To protect the package from unnecessary collisions, it leverages the obstacle-avoidance locomotion skill when moving to the bell and before pushing the package. We test replan-on-interruption as shown in Fig. 5. During the execution of the delivery task, after successfully pressing the doorbell with no response, we manually interrupt the robot’s policy execution and allow it to press the button to open the door. The robot is able to replan based on the new conditions, integrating the previous plan and generating new code to complete the task.
Limitations: Currently, the reasoning layer operates over a fixed set of low-level robot skills crafted by experts. Although the system employs a closed-loop structure with the replanner to handle unexpected events, it remains limited to relatively simple scenarios and has yet to be thoroughly evaluated in highly complex or unstructured failure cases.
@@ V Conclusion
In summary, we present an LLM-based system for tackling long-horizon tasks with a quadruped robot that extends the autonomy of quadrupeds from shorter-term motions for specific skills to long-range complex behaviors that require orchestrating multiple locomotion and manipulation policies. The high-level task planning is enabled with our design of multiple LLM agents for decomposing the whole task into robot actions and then grounded to parameterized robot API calls. The low-level control that exploits the agility of quadrupeds to unlock rich environment interactions is instantiated with model-free RL. Building on this work, further research could extend this closed-loop approach to handle more complex tasks and discover new skills autonomously when confronted with unprecedented circumstances.